OpenStudy (anonymous):

A hybrid car can travel 40 miles per gallon. Approximately how many gallons of fuel will the car need to travel 20 km? [1 mile = 1.6 km] a.) 0.3 b.) 0.8 c.) 2.5 d.) 3.2

OpenStudy (mathmale):

You may find it interesting and helpful to watch your units of measurement carefully. We want to know the number of gallons of gas needed to travel a certain distance, when the fuel economy (miles per gallon) and the distance to be traveled are given. Fuel econ:\[\frac{ 40 mi }{ 1gal }\]and distance: 20 km.

OpenStudy (mathmale):

It'd be easier to convert 20 km into miles than to convert 40 mi / 1 gal into Km / liter (metric system).

OpenStudy (mathmale):

So, let's keep \[\frac{ 40 mi }{ 1gal }\] as the fuel economy rating of the vehicle, anc convert 20 km into miles.

OpenStudy (mathmale):

1 mile is approx. 1.6 Km, and 1 Km is approx. 0.621 mile. Deriving a km->miles conversion factor from the first identity will enable us to eliminate the unit of measurement "Km" in favor of "miles." \[If 1mile=1.6 km, \]then the appropriate conversion factor is \[\frac{ 1mile }{ 1.6 km }.\]

OpenStudy (mathmale):

Application: convert 20Km into miles. Multiply 20 Km by \[\frac{ 1mile }{ 1.6 km }.\]

OpenStudy (mathmale):

We get the following:\[\frac{ 1mile }{ 1.6Km }*20 Km\]

OpenStudy (mathmale):

The unit of measurement Km cancels from numerator and denominator, leaving \[\frac{ 20 miles }{ 1.6}\approx 12.5 miles\]

OpenStudy (mathmale):

Note that dividing the distance, 12.5 miles, by the fuel economy, \[\frac{ 40 miles }{ 1gallon}\]

OpenStudy (mathmale):

produces a quantity with units "gallons," as required. Thus, we calculate\[\frac{ 12.5miles }{ \frac{ 40 miles }{ gallon } }=\frac{ 12.5 }{ 40 }gallons\]

OpenStudy (mathmale):

Reduce that to lowest terms. The answer will be a fraction of a gallon.

OpenStudy (anonymous):

12.5/40 = 0.3125

OpenStudy (anonymous):

@mathmale

OpenStudy (mathmale):

Neffy: 0.3125 what? Important to include units of measurement.

OpenStudy (anonymous):

@mathmale 0.3 gallons. So that is the answer?

OpenStudy (mathmale):

Yes, very good. That's "approximately" how much fuel would be needed.

OpenStudy (anonymous):

@mathmale okay thank you!

OpenStudy (mathmale):

My great pleasure! Good luck and Happy New Year!

OpenStudy (anonymous):

@mathmale You too!